OpenStudy (anonymous):

Solve 5c + 4 = -26 Answer -6 3 6 -4.4

OpenStudy (anonymous):

c=-6

OpenStudy (anonymous):

-6

OpenStudy (anonymous):

a

OpenStudy (anonymous):

5c+4 = -26 => 5c=-26-4=-30 =>c=-30/5 =>c=-6

OpenStudy (ars22):

i also have -6